Thank you so much for your help with this. My brain stopped and I need a push. Can you bring me through this problem please...
"Use a quadratic equation to determine the two numbers that satisfy the situation: Their sum is 9 and their product is 20."?

Answer by venugopalramana(3286)

Thank you so much for your help with this. My brain stopped and I need a push. Can you bring me through this problem please...
"Use a quadratic equation to determine the two numbers -SAY X AND Y
that satisfy the situation: Their sum is 9
X+Y=9..............................................I
and their product is 20."?
XY=20..................................................II
(X+Y)^2-(X-Y)^2=4XY
9^2-(X-Y)^2=4*20=80
(X-Y)^2=81-80=1
X-Y=1....................................III
EQN.II+EQN.III
2X=10
X=5
Y=9-X=9-5=4
SO THE 2 NUMBERS ARE 5 AND 4